OpenStudy (jennychan12):

How would you do integral (cosx/(ln(x+1))dx?

OpenStudy (anonymous):

Just for anyone else, the integral is: \[\int\limits \frac{ \cos x }{ \ln(x+1) } dx\]

OpenStudy (jennychan12):

\[\int\limits \frac{ \cos x }{ \ln(x+1) }dx\]

OpenStudy (anonymous):

Moving onto actually solving it.

OpenStudy (jennychan12):

-_- i'm too slow just fyi, no trig sub or integration by parts

zepdrix (zepdrix):

Hmm then how are you suppose to approach this one? Wolfram didn't come up with a solution. Does this involve Power Series or something? :o

OpenStudy (anonymous):

@zepdrix, it is possible, let me figure it out.

OpenStudy (anonymous):

But do you know?

zepdrix (zepdrix):

@Zelda , please erase your previous comment, and don't talk so vulgar on here... :(

zepdrix (zepdrix):

No I don't know how to do this one I'm afraid :c

OpenStudy (anonymous):

Sorry, just got frustrated, noticed that 3/4ths in.

OpenStudy (anonymous):

I'm fairly certain there's no closed form solution.

OpenStudy (anonymous):

It's actually possible. But requires a power series. @jennychan12, is that okay? Using a power series?

OpenStudy (jennychan12):

what is that? sorry. i'm not good with names. but i think i learned it? is it the same as power rule or something like that?

OpenStudy (anonymous):

No, not a power rule. http://en.wikipedia.org/wiki/Power_series BTW, what level calc are you in?

OpenStudy (jennychan12):

calc Ab

OpenStudy (anonymous):

So I'm assuming you haven't learned the taylor series then?

OpenStudy (jennychan12):

no... -_- what. my teacher gave me ANOTHER BC question. er...something i haven't learned yet. *sigh* my teacher should really make her own practice tests...

OpenStudy (anonymous):

Yea, she did. Would you still like to know how? Also, this actually goes into Calc III a bit, not even BC fully explores series.

OpenStudy (anonymous):

Not sure, but I highly doubt it has an elementary anti-derivative.

OpenStudy (jennychan12):

uhh. i would, but i have a quiz on monday and i don't wanna learn new stuff that my teacher hasn't really taught yet.

OpenStudy (anonymous):

@wio, it actually doesn't. You need a power series to integrate it.

OpenStudy (anonymous):

@jennychan12, in that case I won't go over it. Just say it's impossible since it has no elementary anti-derivative

OpenStudy (jennychan12):

ok haha ok.

OpenStudy (jennychan12):

thanks for your help tho! @Zelda :)

OpenStudy (anonymous):

@jennychan12, no prob. If you want it done later I can show you how.

OpenStudy (jennychan12):

nahh, not now. my brain might explode. haha :/